# Gemstones: A Model Suite for Multi-Faceted Scaling Laws

### v1

@@ 4.3 The Price of Stepping Off the Scaling Law
Industry models often use fewer parameters and train on more tokens than prescribed in prior work (see Figure 5). We find the impact of overtraining a smaller model on predicted loss to be small. Combining this with Figure 7, where wider models are predicted to be optimal in terms of GPU hours, reinforces the message that FLOPs optimality is not the end of the story for training models. Trading some FLOPs optimality for time optimality necessarily means over-/under-training, but Figure 8 suggests the difference is marginal. We believe this combined evidence makes significant progress towards explaining the differences between the prescriptions found in prior work and training choices observed in the wild out in industry.
5 Limitations and Conclusions
We hope this work encourages a rich range of future work on the impact of width and depth within modern transformer architectures using the large amount of open source artifacts we produce. Future work should also extend to other hyperparameters involved in training transformer architectures, such as the expansion factor. Although we endeavor to make our laws as generalizable as possible, we still expect that their applicability declines in training set-ups very different from our own.
Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Software and Data
We train all models using a fork of litgpt (AI, 2023) enhanced with AxoNN (Singh & Bhatele, 2022; Singh et al., 2024) tensor parallelism. We open source all models used in our analysis to Hugging Face (Wolf et al., 2020) and the logging from training on Weights and Biases in json format. All scaling-law fitting code is released on GitHub, with training code to be released shortly after publication.
Appendix B Extended Related Works
Scaling laws are broadly applied to many areas of machine learning, such as machine translation. Ghorbani et al. (2021) split the parameters term into two, one for each of encoder and decoder, and similarly to Gordon et al. (2021) analyze the relationship between BLEU scores and scaling laws. Zhang et al. (2022) and Bansal et al. (2022) study the impact of architecture choice on the scaling law, finding increasing data or parameters can compensate for worse architectural decisions.
Scaling laws have also been applied to sparse architectures. Clark et al. (2022) analyze how the number of experts can be used in the law, studying both linear and quadratic interactions for many types of routing models. Yun et al. (2024) extend this, analyzing the trade offs between optimal training and optimal inference. Krajewski et al. (2024) find that with optimal settings a Mixture of Experts model always outperforms a transformer model at any computational budget. Frantar et al. (2023) focus on weight sparsity within foundation models, adding a multiplicative parameter on the parameters term in the law.
These techniques are not limited to generative text modeling only; they have also been applied to multi-model models. Henighan et al. (2020) find optimal model size can be described as a power law for model modeling including images and video. The authors also find that model size does not help ‘strong generalization’ for problem solving. Aghajanyan et al. (2023) analyze text, images, code and speech, presenting a scaling law to describe the competition between these modalities and describe a regime for optimal hyperparameter transfer from the unimodal to multimodal regimes. Liang et al. (2024) look at scaling laws for diffusion transformer models. Li et al. (2024a) analyze scaling laws for vision encoder commonly used to encode image inputs for transformer model backbones, finding increasing the size of the encoder alone can lead to performance degradation in some cases.
Further analyses using scaling laws have extended to analyzing finetuning and data limited scaling. Hernandez et al. (2021) find that finetuning is much more compute efficient when the pretraining ignored. Zhang et al. (2024) study parameter efficient finetuning regimes find a multiplicative law is better for the finetuning setting than the classical additive law used by others. Muennighoff et al. (2023) analyze the data constrained training regimes, finding epoching data up to four times is as good as training on deduplicated data in terms of reducing loss.
Appendix C Ablations for Approach 1
C.1 Extended Paper Figures
In Figure 9, we plot an extended version of the Approach 1 plot we present in Figure 4.
Figure 9: Extended Approach 1 plot from Figure 4, including tokens and parameters axes. As in Figure 4, we present an analysis over FLOPs on the left and over GPU hours take to train on the right.
C.2 Alternative Learning Rates
In Figure 10, we present the Approach 1 prescription when fitting on the learning rate ablation data.
Figure 10: Approach 1 fitted on the learning rate ablation dataset. As in Figure 4, we present an analysis over FLOPs on the left and over GPU hours take to train on the right.
C.3 Cooldown
In Figure 11, we present the Approach 1 prescription when fitting on the cooldown ablation data.
Figure 11: Approach 1 fitted on the cooldown ablation dataset. As in Figure 4, we present an analysis over FLOPs on the left and over GPU hours take to train on the right.
Appendix D Ablations for Approach 3
D.1 Extended Paper Figures
In Figure 12, we plot an extended version of the Approach 3 plot we present in Figure 5.
Figure 12: Extended Approach 3 plot from Figure 5, including tokens, parameters, width and depth axes. This also includes a comparison of equation 4 that Hoffmann et al. (2022) propose and a brute force plotting method which includes our FLOPs counting for laws of the form seen in Equation 1. We show these are blue and red lines respectively in the right column seeing only a minor difference in the outcome. We remark that industry models fall systematically below our parameter per FLOPs prescriptions, or equivalently above our token per FLOPs prescriptions.
D.2 Alternative Learning Rates
In Figure 13, fit Approach 3 laws to the dataset we trained with half of the optimal learning rate as described in Section 3.1. We again use a brute force approach as above to plot the results but to allow for precise comparison with later ablations we ignore the integer constraints on the number of heads, still enforcing that at least one head number be in every model. We remove this constraint for all models shown in Section D.2, Section D.3, Section D.4 and Section D.5.
Figure 13: Approach 3 fitted on the learning rate ablation dataset. As in Figure 12, we present an analysis over laws of the from shown in Equation 2 on the left and laws of the form shown in Equation 1 on the right.
D.3 Cooldown
In Figure 14, we fit Approach 3 laws to the subset of data for models for which we linearly decreased the learning rate to zero for.
Figure 14: Approach 3 fitted on the cooldown ablation dataset. As in Figure 12, we present an analysis over laws of the from shown in Equation 2 on the left and laws of the form shown in Equation 1 on the right.
D.4 Removing Smaller Token Counts
In Figure 15, we present the Approach 3 prescription when fitting on a dataset where all token counts less than $120$ billion are removed.
Figure 15: Approach 3 fitted on a dataset where all token counts less than $120$ billion are removed. As in Figure 12, we present an analysis over laws of the from shown in Equation 2 on the left and laws of the form shown in Equation 1 on the right.
D.5 Varying Delta in the Huber loss
So far we have fit all approach three laws with a Huber loss delta of $10^{-4}$. We now ablate this decision by refitting all laws with a delta of $10^{-3}$. We use an extremely large grid search of over $4$ million initializations for the width-depth based law when fitting.
To begin we show the prescriptions of the Approach 3 laws if the integer constraints are removed, as we did for the learning rate and cooldown ablations in Figures 13 and 14 respectively.
Figure 16: Following from Figure 5, this plot removes the integer constraints when optimizing. As in Figure 12, we present an analysis over laws of the from shown in Equation 2 on the left and laws of the form shown in Equation 1 on the right.
We now compare all Approach 3 laws found with the increased delta. Specifically, we plot the full dataset laws with delta of $10^{-4}$ in Figure 16 and with $10^{-3}$ in Figure 17. We plot the learning rate ablation laws with delta of $10^{-4}$ in Figure 13 and with $10^{-3}$ in Figure 18. We plot the cooldown ablation laws with delta of $10^{-4}$ in Figure 14 and with $10^{-3}$ in Figure 19. In these figures, we see the difference for the full dataset, cooldown and learning rate ablations laws is minimal when changing the delta. We conclude with a cautionary figure about size of the grid search and the delta used in the Huber loss. In Figure 20, where we plot the exponents found by optimizing the Huber loss versus the size of the grid search used for optimization. We see that a delta of $10^{-5}$ is unstable for smaller grid sizes and including more tokens in the fitting data generally increases stability of the exponents found during optimization.
Figure 17: Approach 3 fitted on our main dataset using $\delta=10^{-3}$ in the Huber loss. Corresponds to Figure 16. As in Figure 12, we present an analysis over laws of the from shown in Equation 2 on the left and laws of the form shown in Equation 1 on the right.
Figure 18: Approach 3 fitted on the learning rate ablation dataset using $\delta=10^{-3}$ in the Huber loss. Corresponds to Figure 13. As in Figure 12, we present an analysis over laws of the from shown in Equation 2 on the left and laws of the form shown in Equation 1 on the right.
Figure 19: Approach 3 fitted on the cooldown ablation dataset using $\delta=10^{-3}$ in the Huber loss. Corresponds to Figure 19. As in Figure 12, we present an analysis over laws of the from shown in Equation 2 on the left and laws of the form shown in Equation 1 on the right.
D.6 Alternative Law Forms
We also experiment with laws of the form shown in Equation 3. In Figure 21, we see that the prescriptions of this law are approximately in line with those of Kaplan et al. (2020). Unlike the laws for shown in Equation 2, these laws tend to prescribe that the width-depth ratio should go to zero as the FLOPs budget increases, i.e. prescribing an infinite-depth model in the infinite-budget limit.
$L(w,d,p,T)=\frac{A}{w^{\alpha}}+\frac{B}{d^{\beta}}+\frac{C}{T^{\gamma}}+\varepsilon$ (3)
Figure 21: Approach 3 fitted on our main dataset using an ansatz of the form shown in Equation 3.
Appendix E Data Sampling
We plot the entire space of all possible models subject to our design constraints discussed in Figure 22. While exploring the impact of finer grained depth differences during our experiments, we decided to add two additional models slightly outside the ±$\pm 5\%$ tolerance band at the $100$M scale; for $width=512$, in addition to the originally chosen depths of $12$ and $13$, we added $11$ and $14$; these appear as a dense collection of $4$ points at the same $width$.
Figure 22: All possible model shapes we could have chosen based on our architecture within ±$\pm 5\%$ are shown as circles. The points we selected are highlighted as stars, including the two extra points we select to have four models of width $512$.
Appendix F Training
@@ F.1 Loss Curves
F.2 Additional Training Complications
@@
We modified the initialization rules and learning rate scaling factors to rescale the depth and layer indices of the model such that if $width/depth>256$ scale variances and learning rates as if the depth of the model was actually $depth^{\prime}=\lceil(width/100)\rceil$. The overall effect of the patch is to initialize and scale learning rates more conservatively, as if the aspect ratio were only $100$ while keeping the original $width$ of the model. We found this allowed us to complete training for a full set of $22$ models out to $350$B tokens for even our most extreme models.
However, after $350B$ tokens, despite these efforts we observed that most extreme models which were patched still diverged anyway. While a partial cause of this could be the constant learning rate scheduler employed during training, concurrent work, from the authors of the original OLMo paper and codebase (Groeneveld et al., 2024) from which we derived some of our choices, reported that the initialization scheme dubbed the “Mitchell-init” is indeed systematically prone to instabilities later on in training (OLMo et al., 2024). While an unfortunate finding, we were unable to rerun all of our experiments due to the consumption of significant non-fungible compute resources in the original experiments.
@@
Our cooldown ablation is from initial experiments below $100B$ tokens of training which do not use the patched learning rates scaling rules. This means there are minor discrepancies between the cooldown ablation and main set of training runs for the widest models from the three largest parameter count groups ($1792\times 7$, $2560\times 8$, $3072\times 12$). We also do not cool down the $100B$ token checkpoint for the $3072\times 12$ model as it was experiencing a loss spike at that final point. Finally, we do not include ablations for the two width $512$ models which do not fall into the ±$\pm 5\%$ boundary of the $100M$ parameter count ($512\times 11$, $512\times 14$) as they were only added to the collection in later experiments.
Appendix G FLOP counting matters
In Figure 24 we show that the common approximation of FLOPs per token$=6\times parameters$, miscounts the true FLOPS by a significant amount.
@@ Appendix G FLOP counting matters
{minted}
python VOCAB_OURS = 50304 SEQ_LEN = 2048 WORLD_BATCH_SIZE = 2048.0 HEAD_SIZE = 128 EXPAND_FACTOR = 4.0
def flops_per_token_gqa( width: NDArray[number] — number, depth: NDArray[number] — number, vocab_size=VOCAB_OURS, queries_per_group=2, seq_len=SEQ_LEN, ): ””” Some details (negligible even for extremely wide models) omitted, including: * numerically stable softmax * softmax addition only being over rows * dot products being only n-1 additions (fused multiply-add exists anyway) ””” num_qheads = width / HEAD_SIZE num_kvheads = ( 2 * num_qheads / queries_per_group )
embeddings = 0 # 0 if sparse lookup, backward FLOPs negligible
attention = 2.0 * seq_len * (num_qheads + num_kvheads) * width * HEAD_SIZE attention += ( 3.5 * seq_len * (num_qheads + num_kvheads / 2) * HEAD_SIZE ) # RoPE, as implemented here/GPT-NeoX # score FLOPs are halved because causal =¿ triangular mask =¿ usable sparsity kq_logits = 1.0 * seq_len * seq_len * HEAD_SIZE * num_qheads softmax = 3.0 * seq_len * seq_len * num_qheads softmax_q_red = 2.0 * seq_len * seq_len * HEAD_SIZE * num_qheads final_linear = 2.0 * seq_len * width * HEAD_SIZE * num_qheads attn_bwd = ( 2.0 * attention + 2.5 * (kq_logits + softmax + softmax_q_red) + 2.0 * final_linear ) * depth attention += kq_logits + softmax + softmax_q_red + final_linear
ffw_size = EXPAND_FACTOR * width dense_block = ( 6.0 * seq_len * width * ffw_size ) # three matmuls instead of usual two because of GEGLU dense_block += ( 10 * seq_len * ffw_size ) # 7 for other ops: 3 for cubic, two additions, two scalar mults dense_block += 2.0 * width * seq_len # both/sandwich residual additions rmsnorm = 2 * 7.0 * width * seq_len
final_rms_norm = 7.0 * width * seq_len # one last RMSNorm final_logits = 2.0 * seq_len * width * vocab_size nonattn_bwd = 2.0 * ( embeddings + depth * (dense_block + rmsnorm) + final_rms_norm + final_logits ) forward_pass = ( embeddings + depth * (attention + dense_block + rmsnorm) + final_rms_norm + final_logits ) backward_pass = attn_bwd + nonattn_bwd # flash attention
return (forward_pass + backward_pass) / seq_len

### v2

@@ Appendix H The Price of Stepping Off the Scaling Law
Industry models often use fewer parameters and train on more tokens than prescribed in prior work. We find the impact of overtraining a smaller model on predicted loss to be small. Combining this with Figure˜16, where wider models are predicted to be optimal in terms of GPU hours, reinforces the message that FLOPs optimality is not the end of the story for training models. Trading some FLOPs optimality for time optimality necessarily means overtraining, but Figure˜17 suggests the difference is marginal. We believe this combined evidence makes significant progress towards explaining the differences between the prescriptions found in prior work and training choices observed in industry.
Figure 17: Quantifying the cost of overtraining. We simulate deviations from our prescriptions to assess their impact on model performance by increasing the optimal token count prescribed by an overtraining factor. We then optimize the model shape to achieve the lowest loss possible at each FLOP budget and overtraining factor. Note that $10^{0}$, or $1\times$, is the prescribed optimal point. We take four FLOP budgets (title of each plot) and plot the loss as a function of overtraining factor and see that under or overtraining increases predicted loss but by only a small amount. We plot the losses of selected open source models on our validation set to help ground the y-axis ranges.
Appendix J Training
@@
J.2 Additional Training Complications
@@
We modified the initialization rules and learning rate scaling factors to rescale the depth and layer indices of the model such that if $width/depth>256$ scale variances and learning rates as if the depth of the model was actually $depth^{\prime}=\lceil(width/100)\rceil$. The overall effect of the patch is to initialize and scale learning rates more conservatively, as if the aspect ratio were only $100$ while keeping the original widthwidth of the model. We found this allowed us to complete training for a full set of $22$ models out to 350350B tokens for even our most extreme models.
However, after $350B$ tokens, despite these efforts we observed that most extreme models which were patched still diverged anyway. While a partial cause of this could be the constant learning rate scheduler employed during training, concurrent work, from the authors of the original OLMo paper and codebase [Groeneveld et al., 2024] from which we derived some of our choices, reported that the initialization scheme dubbed the “Mitchell-init” is indeed systematically prone to instabilities later on in training [OLMo et al., 2024]. While an unfortunate finding, we were unable to rerun all of our experiments due to the consumption of significant non-fungible compute resources in the original experiments.
@@
Our cooldown ablation is from initial experiments below $100B$ tokens of training which do not use the patched learning rates scaling rules. This means there are minor discrepancies between the cooldown ablation and main set of training runs for the widest models from the three largest parameter count groups ($1792\times 7$, $2560\times 8$, $3072\times 12$). We also do not cool down the $100B$ token checkpoint for the $3072\times 12$ model, as it was experiencing a loss spike at that final point. Finally, we do not include ablations for the two width $512$ models which do not fall into the $\pm 5\%$ boundary of the $100M$ parameter count ($512\times 11$, $512\times 14$), as they were only added to the collection in later experiments.
Appendix K Ablations for Approach 1
K.1 Extended Paper Figures
In Figure˜19, we plot an extended version of the Approach 1 plot we present in Figure˜3.
Figure 19: Extended Approach 1 plot from Figure˜3, including tokens and parameters axes. As in Figure˜3, we present an analysis over FLOPs on the left and over GPU hours take to train on the right.
K.2 Alternative Learning Rates
In Figure˜20, we present the Approach 1 prescription when fitting on the learning rate ablation data.
Figure 20: Approach 1 fitted on the learning rate ablation dataset. As in Figure˜3, we present an analysis over FLOPs on the left and over GPU hours take to train on the right.
K.3 Cooldown
In Figure˜21, we present the Approach 1 prescription when fitting on the cooldown ablation data.
Figure 21: Approach 1 fitted on the cooldown ablation dataset. As in Figure˜3, we present an analysis over FLOPs on the left and over GPU hours take to train on the right.
K.4 Varying Delta in the Huber loss
In Figure˜22, where we plot the exponents found by optimizing the Huber loss versus the size of the grid search used for optimization. We see that a delta of $10^{-5}$ is unstable for smaller grid sizes and including more tokens in the fitting data generally increases stability of the exponents found during optimization.
Appendix L FLOP counting matters
[…]
@@
VOCAB_OURS = 50304
SEQ_LEN = 2048
WORLD_BATCH_SIZE = 2048.0
HEAD_SIZE = 128
EXPAND_FACTOR = 4.0
[…]
width: NDArray[number] | number,
depth: NDArray[number] | number,
vocab_size=VOCAB_OURS,
queries_per_group=2,
seq_len=SEQ_LEN,
Some details (negligible even for extremely wide models) omitted, including:
* numerically stable softmax
* softmax addition only being over rows
* dot products being only n-1 additions (fused multiply-add exists anyway)
num_qheads = width / HEAD_SIZE
num_kvheads = (
2 * num_qheads / queries_per_group
[…]
attention += (
3.5 * seq_len * (num_qheads + num_kvheads / 2) * HEAD_SIZE
) # RoPE, as implemented here/GPT-NeoX
# score FLOPs are halved because causal => triangular mask => usable sparsity
kq_logits = 1.0 * seq_len * seq_len * HEAD_SIZE * num_qheads
softmax = 3.0 * seq_len * seq_len * num_qheads
softmax_q_red = 2.0 * seq_len * seq_len * HEAD_SIZE * num_qheads
final_linear = 2.0 * seq_len * width * HEAD_SIZE * num_qheads
attn_bwd = (
2.0 * attention
+ 2.5 * (kq_logits + softmax + softmax_q_red)
+ 2.0 * final_linear
) * depth
attention += kq_logits + softmax + softmax_q_red + final_linear
[…]
dense_block = (
6.0 * seq_len * width * ffw_size
) # three matmuls instead of usual two because of GEGLU
dense_block += (
10 * seq_len * ffw_size
) # 7 for other ops: 3 for cubic, two additions, two scalar mults
dense_block += 2.0 * width * seq_len # both/sandwich residual additions
rmsnorm = 2 * 7.0 * width * seq_len
[…]
final_logits = 2.0 * seq_len * width * vocab_size
nonattn_bwd = 2.0 * (
embeddings + depth * (dense_block + rmsnorm) + final_rms_norm + final_logits
forward_pass = (
embeddings
+ depth * (attention + dense_block + rmsnorm)
+ final_rms_norm
+ final_logits
backward_pass = attn_bwd + nonattn_bwd # flash attention
[…]